OpenStudy (anonymous):

If it requires 18.2 milliliters of 0.45 molar barium hydroxide to neutralize 38.5 milliliters of nitric acid, solve for the molarity of the nitric acid solution. Show all of the work used to solve this problem. Unbalanced equation: Ba(OH)2 + HNO3 yields Ba(NO3)2 + H2O

OpenStudy (anonymous):

I balanced the equation already, Ba(OH)2 + 2 HNO3 = Ba(NO3)2 + 2H2O, can somebody please help me from here?

OpenStudy (anonymous):

@abb0t @Frostbite @primeralph @tcarroll010 @thomaster @uri @zzr0ck3r can one of you please help me learn how to do this? I would really appreciate it!

OpenStudy (primeralph):

Hold on.

OpenStudy (anonymous):

@primeralph would you mind typing the numbers out and explaining them so i can comprehend a little bit better please?

OpenStudy (primeralph):

M = n/V, M = molarity, n= number of moles, V = volume n Ratio is 1:2 So, (M_1*V_1)/(M_2*V_2) = 1/2 (18.2*0.45)/(x*38.5) = 1/2

OpenStudy (abb0t):

molarity is moles/ volume. So find the moles and then divide by volume.

OpenStudy (primeralph):